OpenStudy (anonymous):

Find the value of k so that the binomial x-1 is a factor of -3x^4+kx^3+6x^2-9x+3

OpenStudy (anonymous):

replace \(x\) by 1, set the result equal to zero, solve for \(k\)

OpenStudy (anonymous):

if \(x-1\) is a factor of \(f(x)=-3x^4+kx^3+6x^2-9x+3\) then \(f(1)=0\) this is because if \(x-1\) is a factor then \(-3x^4+kx^3+6x^2-9x+3=(x-1)(something)\) and so \(x=1\) gets you \((1-1)(something )=0\) therefore \[-3+k+6-9+3=0\] and you can find \(k\) from that equation

OpenStudy (anonymous):

Thanks :)